In a wire, when the elongation is 2 cm, the energy stored is E. if the wire is stretched by 10 cm, then the energy stored in the wire will be

(a) E (b) 5E

(c) 25E (d) 25/2E

Grade:11

1 Answers

Vikas TU
3 years ago

We know, potential energy of spring is given by (1/2)kx² where k is spring constant and x is stretched distance.

Here , x=2cm so, U=(1/2)k*2²=2k

For, x=10cm :-

E=(1/2)k*10²=50k=25(2k)=25U.

Final potential energy will be 25U.